OpenStudy (anonymous):

f the circle x^2 - 4x + y^2 + 2y = 4 is translated 3 units to the right and 1 unit down, what is the center of the circle? A) (2, 0) B) (2, -1) C) (5, -2) D) (-2, 1)

OpenStudy (anonymous):

@PercyJackson&theΩ

OpenStudy (astrophysics):

Complete the square

OpenStudy (anonymous):

can u draw it out?

OpenStudy (astrophysics):

It will then be in \[(x-h)^2+(y-k)^2=r^2\] where (h,k) is the center

OpenStudy (anonymous):

to complete the suare do u have to use b/2a?

OpenStudy (anonymous):

square*

OpenStudy (astrophysics):

Yea you can

OpenStudy (astrophysics):

But I doubt you'll get anything out of it, just complete the square.

OpenStudy (anonymous):

theoretically would he not have to do x^2-6x=x^2-6x+9-9 this is just an example

OpenStudy (astrophysics):

Just show me what you're thinking, do it as if I wasn't here I guess.

OpenStudy (anonymous):